# Inversion of Magnetic Data using Learned Dictionaries and Scale Space

### v1

@@ 4.2.3 Learned ${\bm{\mathrm{\Psi}}}$:
Both the shared ${\bm{\mathrm{\Psi}}}$ and the unrolled ${\bm{\mathrm{\Psi}}}^{(i)}$ used a 3D convolution of kernel dimensions $7\times 7\times 7$ with a channel dimension $C=9$ and a total of $16$ iterations for direct comparison. The unrolled dictionary had a new set of weights on each iteration adding complexity compared to the single learned dictionary of shared weights.
@@ 4.3 Comparison of model recovery
A comparison of the normalized mean square error (nMSE) between the four different methods is shown in Table 2. Qualitative comparison between different methods for the recovery of 3D models has been shown for some representative test set samples in Fig. 2. In addition, we compare the distribution of losses between variational methods and unrolled ${\bm{\mathrm{\Psi}}}$ in Fig. 1 (left) and between shared ${\bm{\mathrm{\Psi}}}$ and unrolled ${\bm{\mathrm{\Psi}}}$ in Fig. 1 (right).
Table 2: Normalized Mean Squared Error (nMSE) for model and data recovery using different methods. Bold and red values indicate the best and the second best method for each recovery type.
@@
Figure 2: Three sample models (01, 11, and 19) from the test set illustrating the quality improvement of unrolled dictionary learning for inversion compared to other methods. Each image highlights different features of the model reconstruction quality achieved using the proposed method.
The results clearly demonstrate a remarkable improvement of learnable dictionaries over fixed regularization or dictionaries. The improvement is not marginal. We observe an over $60\%$ improvement in the recovery on average compared to variational approaches and over $55\%$ compared to fixed dictionary. Comparing the unrolled dictionary that represents a scale space approach to the learned single dictionary we observe that the learned variable dictionary is significantly better, emphasizing the advantage of scale space methods over variational methods.
5 Summary and conclusion

### v2

@@ 4.2.3 Learned ${\bm{\mathrm{\Psi}}}$:
Both the shared ${\bm{\mathrm{\Psi}}}$ and the unrolled ${\bm{\mathrm{\Psi}}}^{(i)}$ used a 3D convolution of kernel dimensions $7\times 7\times 7$ with a channel dimension $C=9$ and a fixed total of $N=16$ iterations in Algorithms 1 and 2 for direct comparison. The unrolled dictionary had a new set of weights on each iteration adding complexity compared to the single learned dictionary of shared weights.
@@ 4.3 Comparison of Model Recovery
A comparison of the normalized mean squared error (nMSE) between the four different methods is shown in Table 2. Qualitative comparison between different methods for the recovery of 3D models has been shown for some representative test set samples in Fig. 2. In addition, we compare the distribution of losses between variational methods and unrolled ${\bm{\mathrm{\Psi}}}$ in Fig. 1 (left) and between shared ${\bm{\mathrm{\Psi}}}$ and unrolled ${\bm{\mathrm{\Psi}}}$ in Fig. 1 (right).
Table 2: Normalized Mean Squared Error (nMSE) for model and data recovery using different methods. Bold and italicized values indicate the best and the second best method for each recovery type.
@@
Figure 2: Three sample models (01, 11, and 19) from the test set illustrating the quality improvement of unrolled dictionary learning for inversion compared to other methods. Each image highlights different features of the model reconstruction quality achieved using the proposed method. The 2D magnetic data from the 3D true model is inverted back to a 3D construction using four different methods. Views of the reconstruction are shown using contour potentials of the magnetic susceptibility.
The results clearly demonstrate a remarkable improvement of learnable dictionaries over fixed regularization or dictionaries. The improvement is not marginal. We observe an over $60\%$ improvement in the recovery on average compared to variational approaches and over $55\%$ compared to fixed dictionary. Comparing the unrolled dictionary that represents a scale space approach to the learned single dictionary in Figure 2, we observe that the learned variable dictionary produces inverse models that capture features deeper below the subsurface and with closer fidelity to the true model, emphasizing the advantage of scale space methods over variational methods.
5 Conclusion